OpenStudy (anonymous):

Choose the correct first step(s) to solve the system using elimination. 4x + 7y = 6 and 6x + y = 2

OpenStudy (anonymous):

divide both sides by X

hero (hero):

All I can say is, you have at least two other people taking the exact same class you are.

hero (hero):

Katin 95 is one of them

OpenStudy (anonymous):

Lol! I go to K12 online public school.

hero (hero):

I know

OpenStudy (anonymous):

Ok? Lol. Can you help me put what Durviel just said into an equation??

hero (hero):

I wouldn't use durviel's method

hero (hero):

I have my own methods

OpenStudy (anonymous):

Oh, well can you use your method and put it in an equation for me? :)

OpenStudy (anonymous):

everyone has different methods. some work better for others. correct?

hero (hero):

durviel, I'm curious to see your method

hero (hero):

By the way, it says use elimination

OpenStudy (anonymous):

i would find x first. then, when i have that answer, i use it to find y. i was taught that way. but maybe your way does work better.

OpenStudy (anonymous):

I just want an answer lol.

hero (hero):

Originally, you said "divide both sides by x". I was waiting to see how you would use that to solve it.

OpenStudy (anonymous):

how would you do it? i am hear to learn so... yeah

hero (hero):

You first

OpenStudy (anonymous):

He just told you his method, Lol.

OpenStudy (anonymous):

one sec....

hero (hero):

I want to see the method where you divide both sides by x only.

OpenStudy (anonymous):

the way i was taught is you find x first. then when you plug in x, you could find y.

OpenStudy (anonymous):

What was that drivel between Durviel and Hero lol...Thought you were supposed to be helping... I will help. So we have: 4x + 7y = 6 and 6x + y = 2 The correct first step is to multiply the second equation by 7 to get: 42x + 7y = 14 We can then eliminate the 7 by subtracting equation 1 (unchanged) from equation 2. :)